# EmoBench-Reddit: A Hierarchical Benchmark for Evaluating the Emotional Intelligence of Multimodal Large Language Models

### v1

@@ Abstract
With the rapid advancement of Multimodal Large Language Models (MLLMs), they have demonstrated exceptional capabilities across a variety of vision-language tasks. However, current evaluation benchmarks predominantly focus on objective visual question answering or captioning, inadequately assessing the models’ ability to understand complex and subjective human emotions. To bridge this gap, we introduce EmoBench-Reddit, a novel, hierarchical benchmark for multimodal emotion understanding. The dataset comprises 350 meticulously curated samples from the social media platform Reddit, each containing an image, associated user-provided text, and an emotion category (sad, humor, sarcasm, happy) confirmed by user flairs. We designed a hierarchical task framework that progresses from basic perception to advanced cognition, with each data point featuring six multiple-choice questions and one open-ended question of increasing difficulty. Perception tasks evaluate the model’s ability to identify basic visual elements (e.g., colors, objects), while cognition tasks require scene reasoning, intent understanding, and deep empathy integrating textual context. We ensured annotation quality through a combination of AI assistance (Claude 4) and manual verification.
We conducted a comprehensive evaluation of nine leading MLLMs, including GPT-5, Gemini-2.5-pro, and GPT-4o, on EmoBench-Reddit. The results reveal several key characteristics of current models: 1) Top-tier models, represented by GPT-5 and Gemini-2.5-pro, exhibit a leading edge in most tasks, especially showing stable performance on basic perception and elementary cognition. 2) All models perform well on basic perception tasks, but their performance drops significantly on advanced cognition tasks, indicating that the leap from "seeing" to "understanding" remains a formidable challenge. 3) The complexity of emotion types significantly impacts model performance; emotions like "sarcasm," which require deep contextual and cultural understanding, are the biggest hurdle for all models, whereas "happy" and "humor" are relatively easier to process. 4) Models with smaller parameter counts show clear capability bottlenecks when handling complex cognitive tasks. This study not only provides an effective tool for assessing the emotional intelligence of MLLMs but also points to critical directions for future model development, namely, strengthening deep reasoning and contextualized understanding of complex human emotions.
EmoBench-Reddit: A Hierarchical Benchmark for Evaluating the Emotional Intelligence of Multimodal Large Language Models

### v3

# Seeing is Not Understanding: A Benchmark on Perception–Cognition Disparities in Large Language Models

@@ Abstract
With the rapid advancement of Multimodal Large Language Models (MLLMs), they have demonstrated exceptional capabilities across a variety of vision-language tasks. However, current evaluation benchmarks predominantly focus on objective visual question answering or captioning, inadequately assessing the models’ ability to understand complex and subjective human emotions. To bridge this gap, we introduce EmoBench-Reddit, a novel, hierarchical benchmark for multimodal emotion understanding. The dataset comprises 350 meticulously curated samples from the social media platform Reddit, each containing an image, associated user-provided text, and an emotion category (sad, humor, sarcasm, happy) confirmed by user flairs. We designed a hierarchical task framework that progresses from basic perception to advanced cognition, with each data point featuring six multiple-choice questions and one open-ended question of increasing difficulty. Perception tasks evaluate the model’s ability to identify basic visual elements (e.g., colors, objects), while cognition tasks require scene reasoning, intent understanding, and deep empathy integrating textual context. We ensured annotation quality through a combination of AI assistance (Claude 4) and manual verification.We conducted a comprehensive evaluation of nine leading MLLMs, including GPT-5, Gemini-2.5-pro, and GPT-4o, on EmoBench-Reddit.
Seeing is Not Understanding: A Benchmark on Perception–Cognition Disparities in Large Language Models